OpenStudy (anonymous):

Solve equation 5(x-9) = 8(x+6)-x. Solve for X.

OpenStudy (anonymous):

Distribute the 5 to x and -9...and the 8 to x and 6

OpenStudy (anonymous):

Can you please show the work?

OpenStudy (anonymous):

5(x-9) = (5*x)+(5*-9) = 5x-45 Now distribute to the other set